OpenStudy (anonymous):

What is the solution of the following system? -2x-y=0 3x-y=10 possible answers are: 1. infinitely many solutions 2. no solutions 3. (3,8) 4. (-3,-8)

OpenStudy (jdoe0001):

recheck your equations, are there any missing negative signs somewhere, or some other value missing?

OpenStudy (anonymous):

I checked its not missing anything @jdoe0001

OpenStudy (jdoe0001):

well... if you solve the 1st equation for "y", what would you get?

OpenStudy (anonymous):

y=2x? @jdoe0001

OpenStudy (jdoe0001):

well... -2x -y + y = 0 + y => -2x = y

OpenStudy (jdoe0001):

so now we know that y = -2x so let us use that in the 2nd equation then \(\bf y = -2x\qquad thus\\ \quad \\ 3x-y=10\implies 3x-(-2x)=10\implies 3x+2x=10\implies 5x=10\\ \quad \\\implies x =2\\ \quad \\ \textit{using our new found "x" in say the 1st equation, we get}\\ \quad \\ -2x-y = 0\implies -2(2)-y=0\implies -4 = y \)

OpenStudy (jdoe0001):

so as you can see, x = 2, y = -4 so the solution will be ( 2 , -4 ) and is not one of your choices

OpenStudy (anonymous):

well now I don't know what to do because those are my only choices lol oh well. could you help me with a different one? @jdoe0001

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Solve the system by substitution. -4.5x-2y=-12.5 3.25x-y=-0.75 I have to show work step by step but I have no idea how to even start it

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

same thing, you'd solve either equation for "y" once you get "y", you use that equivalent in the other equation by SUBSTITUTING the "y'" there

OpenStudy (jdoe0001):

that's all I did, you solved for "y" on the 1st one, and I used that "y" equivalent in the 2nd one

OpenStudy (anonymous):

could you help me with solving for y on the first equation and then ill do the second one myself because I think im soling for y incorrectly @jdoe0001

OpenStudy (jdoe0001):

okie dokie \(\bf -4.5x-2y=-12.5\qquad \textit{adding 12.5 to both sides}\\ \quad \\ -4.5x-2y+12.5=\cancel{-12.5+12.5}\implies -4.5x-2y+12.5=0\ \quad \\\qquad \textit{adding 2y now}\\ \quad \\ -4.5x\cancel{-2y+2y}+12.5=0+2y\implies -4.5x+12.5=2y\ \quad \\\qquad \textit{dividing by 2}\\ \quad \\ \cfrac{-4.5x+12.5}{2}=\cfrac{\cancel{2}y}{\cancel{2}}\implies \cfrac{-4.5x+12.5}{2}=y\)

OpenStudy (anonymous):

so y in that eqaution= 2.25+6.25? @jdoe0001

OpenStudy (jdoe0001):

-2.25x + 6.25 <--- yes

OpenStudy (anonymous):

which is 4 right so I sub 4 in for y in the second equation?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

4? what do you mean? y = 4?

OpenStudy (anonymous):

yeah 6.25-2.25 is 4? @jdoe0001

OpenStudy (jdoe0001):

\(\Large -2.25\color{red}{x} + 6.25 = y\)

OpenStudy (anonymous):

so 3.25x-2.25x+6.25=-0.75 for the second equation

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, kinda, recall that - times - = + and that - times + = - same sign multiplication gives use POSITIVE different sign multiplication gives use NEGATIVE \(\bf 3.25x-y=-0.75\implies 3.25x-(-2.25x + 6.25)=-0.75\\ \quad \\ \color{blue}{-(-2.25x + 6.25)\implies +2.25x-6.25}\qquad thus\\ \quad \\ 3.25x-(-2.25x + 6.25)=-0.75\implies 3.25x+2.25x-6.25=-0.75\)

OpenStudy (anonymous):

oh okay thanks! @jdoe0001

OpenStudy (jdoe0001):

yw